OpenStudy (anonymous):

how do you solve: e^(x+5)=(e^x)+6 please

OpenStudy (anonymous):

e^(x+5)=(e^x)+6 e^x ( e^5 -1) = 6 e^x = 6/ ( e^5 -1) x= ln [ 6/ (e^5-1)]

OpenStudy (anonymous):

that is correct, thank you! may i ask though how you derived (e^5-1)?

OpenStudy (anonymous):

e^(x+5) = (e^x)*(e^5)

OpenStudy (anonymous):

im so sorry, i still don't understand how you get e^5-1 from that

OpenStudy (anonymous):

e^(x+5)=(e^x)+6 take out the common factor of e^(x+5) and (e^x), which is e^x: e^x ( e^5 -1) = 6 e^x = 6/ ( e^5 -1) x= ln [ 6/ (e^5-1)] got it ??????

OpenStudy (anonymous):

yesss thank you